Write an english version of the following equations;
a) 2x-5=+1 to write this in english I came up with two x is less than 5 plus one more. ( I know this is not right, I need help with these)
b) 2(x+2)=14
c) n+5=n/2-6
d) 7-3a=5+a
I don't know how to translate these into english word problems, help please as you can see I messed up A bad!!!

Answer by jayanthihemkos@yahoo.com(87)

Write an english version of the following equations;
a) 2x-5=+1 to write this in english I came up with two x is less than 5 plus one more. ( I know this is not right, I need help with these)
b) 2(x+2)=14
c) n+5=n/2-6
d) 7-3a=5+a
I don't know how to translate these into english word problems, help please as you can see I messed up A bad!!!
a) 2x-5=+1
Ans: Five less than two times x is equal to one
solution

2x-5 = 1

2x -5 +5 = 1+5
2x = 6
2x/2 = 6/2
x = 3
Hence the solution
b) 2(x+2)=14
Ans: = Twice the sum of x and 2 is equal to fourteen
solution:
2(x+2)=14
2x + 4 = 14 (open the paranthesis)
2x + 4 -4 = 14 -4

2x = 10
2x/2 = 10 /2
x = 5
Hence the solution
c) n+5=n/2-6
ANS: sum of n and five is equal to six less than n divided by two
or
Five more than n is equal to six less than n divided by 2
solution:

n+5=n/2-6
Multiply through out by 2
2n +10 = n-12
2n + 10 -n = n -n -12
n + 10 = -12
n + 10 -10 = -12 -10
n = -22
Hence the solution

d) 7-3a=5+a
Difference of seven and three times a is equal to the sum of five and a
solution:
d) 7-3a=5+a
7 -7 -3a = 5- 7 + a
-3a = -2 + a
- 3a -a = -2 + a - a
- 4a = -2
4a = 2
4a/4 = 2/ 4
a = 1/2
Hence the solution